OpenStudy (anonymous):

Use the graph of y = f(x) to graph the given function g. g(x) = -2f(x) I'm not sure how of how to solve this..

ganeshie8 (ganeshie8):

where is the graph

OpenStudy (anonymous):

There's only graphs for the answer choices http://awesomescreenshot.com/0b632rtfd5 http://awesomescreenshot.com/0b032rtt07

OpenStudy (kmeis002):

One way to approach this problem is to note that g(x) only modifies the range of f(x). That implies the domain will not change. Looking at the graphs, we can then find the correct graph.

OpenStudy (anonymous):

I'm not exactly sure of what you mean by that but in all of the graphs I see a point on (6, ...) for all of them... but the second graph has (6,..) for both lines e.e

OpenStudy (kmeis002):

Well here's an example. Lets say f(x) has these points associated with it: (1,2) (2,4) and (4,6) Then if g(x) = -2(f(x)) that is telling us for any x we just take the output of f(x) and multiply it by -2 so g(1) = -2*f(1) = -2(2) = -4 g(2) = -2*(f(2)) = -2*4 = -8 In this case, g(x) and f(x) domain will remain the same. If the domain of f(x) is [1,4] then g(x) has a domain of [1,4], but their range values differ.

OpenStudy (anonymous):

>.< I'm sorry I'm so confused...I'm trying to apply what you just said to me to thr graphs and their points but I don't really know where I'm supposed to be looking at

OpenStudy (kmeis002):

Ah, alright. So Let me graph an arbitrary f(x) with some specific points. I will use the points listed in my previous post: (1,2), (2,4), and (4,6). The graph could look something like the image I posted here. 

OpenStudy (kmeis002):

Then if we graph g(x), we will get the points (1, -4), (2, -8), and (4, -12) because g(x) takes the outputs of f(x) and multiplies them by -2.

OpenStudy (kmeis002):

 Notice how the points align with each other on the x-axis, this is because g(x) only effected the output of f(x)